# AI-Generated Video Detection via Perceptual Straightening

### v1

@@ Abstract
The rapid advancement of generative AI enables highly realistic synthetic videos, posing significant challenges for content authentication and raising urgent concerns about misuse. Existing detection methods often struggle with generalization and capturing subtle temporal inconsistencies. We propose ReStraV(Representation Straightening for Video), a novel approach to distinguish natural from AI-generated videos. Inspired by the “perceptual straightening” hypothesis [1, 2]—which suggests real-world video trajectories become more straight in neural representation domain—we analyze deviations from this expected geometric property. Using a pre-trained self-supervised vision transformer (DINOv2), we quantify the temporal curvature and stepwise distance in the model’s representation domain. We aggregate statistics of these measures for each video and train a classifier. Our analysis shows that AI-generated videos exhibit significantly different curvature and distance patterns compared to real videos. A lightweight classifier achieves state-of-the-art detection performance (e.g., 97.17% accuracy and 98.63% AUROC on the VidProM benchmark [3]), substantially outperforming existing image- and video-based methods. ReStraV is computationally efficient, it is offering a low-cost and effective detection solution. This work provides new insights into using neural representation geometry for AI-generated video detection.
Figure 1: ReStraV approach to detect AI-generated videos, inspired by the “perceptual straightening” hypothesis. Frames encoded by a self-supervised encoder (SSE); e.g., DINOv2, yield embedding trajectories ($z_{i}$). Consistent with the hypothesis, natural videos tend to form “straighter” trajectories in this representation space compared to AI-generated ones. These geometric distinctions, such as differences in temporal curvature (Eq. 1), provide a discriminative signal for fake video detection.
@@ 1 Introduction
This work explores neural representational distance and curvature (formally defined in Eq. 1) as discriminative signals for fake video detection. An overview of ReStraV is provided in Fig. 1. According to the perceptual straightening hypothesis, natural inputs map to straight paths in neural representations while unnatural sequences form curved trajectories [1, 2]. This has been verified in neuroscience, psychophysics, on CNNs and LLMs [1, 23]. It is motivated by the idea that predictive coding might favor straight temporal trajectories in latent space because they are more predictable.
Taking inspiration, in this work, we hypothesize a distinction between natural and AI generated videos in artificial neural networks (ANNs). While ANNs may not perfectly replicate biological straightening [1, 2, 24], we expect their learned representations to show AI-generated videos as more curved in activation space than real videos. We surmise synthetic videos exhibit curvature patterns deviating from the lower curvature trajectories of real events, supported by differing ANN representational dynamics for natural versus artificial videos [24], as illustrated in Fig. 2A and Fig. 4.
To test this hypothesis, we use the DINOv2 ViT-S/14 pretrained visual encoder [25], chosen for its sensitivity to generative artifacts (Fig. 2B). For each video, we extract frame-level complete set of patch embeddings and Classify token (CLS) from DINOv2’s final transformer block (block.11). From this trajectory, we quantify local curvature (angle between successive displacement vectors, measuring path bending) and stepwise distance (change magnitude between consecutive frame representations), as in Eq. 1. We then derive descriptive statistics (mean, variance, min, max; examples in Fig. 5) from these per-video time series of curvature and distance. These aggregated geometric features (Section 5) are used by a lightweight classifier (Section 6) to distinguish real from AI content.
ReStraV re-purposes DINOv2 as a “feature space” for temporal anomalies. DINOv2’s extensive training on natural data provides a latent space where real video trajectories should be characteristically smooth or “straight” (Fig. 2A, Fig. 4). Deviations, like increased jitter or erratic curvature often in AI videos, become discernible geometric signals of synthetic origin. Importantly, ReStraV is computationally efficient, processing videos in approximately 48 ms end-to-end (including DINOv2 forward pass). ReStraV is thus a low-cost alternative to resource-intensive methods (details in Section 6). By exploiting ANN’s activation dynamics, ReStraV offers a simple, interpretable AI-video detection approach (experimental validation in Section 7). Our contributions are as follows:
@@
Figure 2: (A) In pixel space (left), video trajectory metrics (curvature, distance; see Eq. 1 for details) between natural vs. AI-generated videos show substantial overlap. In contrast, DINOv2 representations (right) straighten natural trajectories, clearly separating natural and AI-generated videos. (B) The mean curvature gap ($\Delta\theta$) between AI-generated and natural videos, evaluated across various visual encoders (optimal layers), shows that curvature is a robust signal across encoders and most pronounced in DINOv2.
@@ 2 Related work: detecting AI–generated videos
Detecting AI-generated video (see Appendix 9 for an overview of generation techniques) is becoming increasingly challenging. Many early detection efforts, including several image-based detectors benchmarked in our work (CNNSpot [27], Fusing [28], Gram-Net [16], FreDect [29], GIA [30], LNP [31], DFD [32], UnivFD [33], focused on spatial or frequency-domain artifacts within individual frames. However, their frame-centric nature limits their efficacy on videos, where temporal consistency is paramount. Such methods may identify low-level statistical deviations but often miss the more holistic temporal unnaturalness.
Dedicated video detectors, such as adapted action recognition models (TSM [34], I3D [35], SlowFast [36]) and Transformer-based architectures (X3D [37], MVIT-V2 [38], VideoSwin [39, 40], TPN [41], UniFormer-V2 [42], TimeSformer [43], aim to learn motion anomalies and temporal inconsistencies. While advancing temporal modeling, these often require extensive training on video data, can be computationally intensive, and may still struggle with generalization across diverse and rapidly evolving AI generative models.
Critically, both image- and video-based approaches that learn specific artifact patterns or complex motion dynamics might overlook more fundamental geometric distortions in the temporal trajectory of neural representations that AI generation can introduce. These geometric properties could offer a more robust and universal signal, as they relate to the intrinsic predictability and smoothness of natural world dynamics, which generative models may not fully capture.
@@ 4 Perceptual straightening of natural videos in DINOv2
We extract $T=24$ frames $\mathcal{Z}=(z_{1},\dots,z_{24})$ by sampling over a $2$-second video duration. This $2$-second window is suitable for the videos considered in Section 7 ($\approx 2-5$s long with $12-30$ FPS), with an temporal based sample frame of $\Delta t=2\text{s}/(24-1)$. Nevertheless, we hypothesize that using longer videos could further enhance performance. Our choice of a $2s$ window with $24$ frames was found to provide an optimal trade-off between high detection accuracy and computational efficiency, as validated in our ablation studies (see 11 for details). Each $x_{i}$ is resized to $224\times 224$ pixels and normalized to $[0,1]$. These preprocessed frames are then encoded by the DINOv2 ViT-S/14 model [25]. The $384$ CLS (Classify) tokens and $196$ patch embedding ($16*16$ patches of the $224*224$ inputs) from its final transformer block (block.11). These token embeddings are then flattened and concatenated to form a single feature vector $z_{i}\in\mathbb{R}^{75648}$. The sequence of these vectors, $\mathcal{Z}$, forms the temporal trajectory in DINOv2’s representation space, from which temporal curvature and distance metrics are computed (Eq. 1).
Although the learned representations of ANNs like DINOv2 may not fully straighten natural video trajectories [24], their activation dynamics exhibit significant and informative differences between natural and AI-generated videos. This distinction is evident when comparing trajectory like curvature and distance (Eq. 1) in pixel versus DINOv2 representation spaces (Fig. 2A). The analysis for Fig. 2A uses natural videos from the Physics-IQ benchmark [26] and synthetic replicas (generated using models like Sora [48], Pika [44], and Runway2 [49]), where each natural-synthetic pair is matched one-to-one in terms of starting scene/objects and duration to isolate generative artifacts. While these geometric metrics show considerable overlap for such matched pairs in pixel space, they differ markedly in DINOv2’s representation space. Consequently, although natural videos are not perfectly linearized in DINOv2 space, their curvature and distance patterns significantly diverge from the irregular trajectories typical of AI-generated content, offering a robust detection signal. Refer to Section 10.1 for trajectories samples.
The DINOv2 ViT-S/14 encoder demonstrates strong differences representing natural versus AI-generated videos. A comparative analysis of eight vision encoders—AlexNet [50], VGG-16 [51], ResNet-50 [52], SIN-ResNet-50 [53], SimCLR-R50 [54], BYOL-R50 [55], CLIP (ViT-B/32) [56], and DINOv2 (ViT-S/14) [25] itself—confirms its superiority (Fig. 2B). Evaluated on paired natural videos from the Physics-IQ benchmark [26] and corresponding synthetic videos, DINOv2 reveals a significantly larger mean curvature difference ($\Delta\theta=\overline{\theta}_{\mathrm{AI}}-\overline{\theta}_{\mathrm{Nat}}$) between their trajectories.
Further evidence of DINOv2’s effectiveness in creating a discriminative representation space is presented in Fig. 4. This figure visualizes t-SNE embeddings of video curvature trajectories from the VideoProM dataset [3], including natural videos and synthetic ones from various generators (such as Pika [44], VideoCraft2 [57], Text2Video-Zero [58], and ModelScope [47]). In raw pixel space (left panel), trajectories from natural and synthetic videos largely overlap. In contrast, within the DINOv2 ViT-S/14 representation space (right panel), natural video trajectories (blue) form a distinct central cluster, while AI-generated trajectories (shades of red) are well-separated and are often grouped by their generative model. This demonstrates that DINOv2’s self-supervised features effectively capture subtle temporal inconsistencies differentiating AI-generated from natural videos, even without specific training on video data or for the detection task itself.
{mdframed}
@@
[ middlelinecolor = black, middlelinewidth = 0.5pt, backgroundcolor = NatBlue!20, roundcorner = 4pt, ] Takeaway 1: By projecting videos into DINOv2’s representation space, geometric trajectory features (curvature & distance, Eq. 1) become indicators of synthetic origin, differentiating AI-generated videos from natural ones in a way that is not possible in raw pixel space.
@@ 5 Analyzing characteristics of perceptual trajectories
We select 50,000 AI-generated samples (10,000 each from Pika [44], VideoCraft2 [59], Text2Video-Zero [58], ModelScope [3], and Sora [60]) from VideoProM [3]. Concurrently, 50,000 natural videos are randomly chosen from DVSC2023 [45]. All videos are DINOv2 (ViT-S/14) encoded, and their aggregated statistical features are computed. Fig. 5 illustrates these aggregated feature distributions. The top row shows distance features ($\mu_{d},\min_{d},\max_{d},\sigma_{d}^{2}$), characterizing inter-frame change magnitude and variability. The bottom row presents corresponding curvature features ($\mu_{\theta},\min_{\theta},\max_{\theta},\sigma_{\theta}^{2}$), reflecting angular changes between consecutive frame transitions.
@@ 6 Video classifier to detect AI-generated content
Given that DINOv2 representation distance $d$ and curvature $\theta$ differs significantly between natural and AI-generated videos, we evaluate if these features can be used in a lightweight, transparent, and easily replicated classifier without raw pixel processing or DINOv2 fine-tuning. We use the dataset from Section 5 and apply a stratified 50/50 train/test split. Class priors are identical, and each subset is balanced among five AI models (Pika [44], VideoCraft2 [59], Text2Video-Zero [58], ModelScope [3] and Sora [60]). We sample frames and we obtain the signals of distance $\{d_{i}\}_{i=1}^{T-1}$ and curvature $\{\theta_{i}^{\circ}\}_{i=1}^{T-2}$ as detailed in Section 5. For classification, we construct a feature vector $y$ per video by combining direct signals and aggregated statistics from these trajectories. Specifically, $y$ concatenates seven distance values $[d_{1},d_{2},\dots,d_{7}]$ and six curvature values: $[\theta_{1}^{\circ},\theta_{2}^{\circ},\dots,\theta_{6}^{\circ}]$; and four statistical descriptors (mean, variance, minimum, maximum) for both $\{d_{i}\}$ and $\{\theta_{i}^{\circ}\}$. This results in a final feature vector $y\in\mathbb{R}^{21}$.
We consider only off-the-shelf models: logistic regression (LR), Gaussian Naive Bayes (GNB), random forest (RF; 400 trees, depth $\leq 6$), gradient boosting (GB; 200 rounds, learning rate 0.1), RBF-kernel SVM (calibrated by Platt scaling), and a two-layer MLP $(64\rightarrow 32)$. We perform no feature engineering or hyperparameter search beyond a 3-fold grid/random sweep. For each classifier, we optimize the decision threshold $\tau^{*}$ on the training set to maximize the F1-score. The chosen threshold $\tau^{*}$ was then applied unchanged to the test set. Inference cost is reported end-to-end ($\text{latency}=T_{\text{DINOv2}}+T_{\text{clf}}$), averaged over the test fold on a single NVIDIA RTX-2080. A DINOv2 forward pass (ViT-S/14, block 11, 8-frame batch) takes 43.6 ms. This constant is added to each classifier time ($T_{\text{clf}}$) in Table 1.
Figure 6: Inference-time ($ms$) vs. accuracy (%) and AUROC (%) for ReStraV’s classifiers.
Table 1: Performance and inference time of ReStraV’s classifiers, balanced 50k/50k natural/AI-generated video test set from VideoProM [3], cf. Section 5 for details. The best scores are bold, second best are underlined and the best method overall is highlighted in blue.
@@
Table 1 summarizes test performance. The MLP achieves the highest accuracy (94.17%), F1-score (96.48%), and AUROC (98.6%) (Fig. 9A), followed by RF. Section 6 visualizes each classifier’s speed/accuracy. Models in the upper-left offer the best cost/benefit. The MLP (highlighted blue in Table 1) achieves top AUROC and F1 while being within $\approx$2ms from GNB. The confusion matrices and ROC (Fig. 9A and B in Section 10.2) confirms low false positive/negative rates (cf. Section 10.2 for decision boundaries visualization).
{mdframed}
[ middlelinecolor = black, middlelinewidth = 0.5pt, backgroundcolor = NatBlue!20, roundcorner = 4pt, ] Takeaway 2: Simple, lightweight classifiers trained directly on geometric trajectory features achieve high classification accuracy ($\approx 94\%$) and AUROC ($\approx 99\%$), offering an efficient detection approach ($\approx 48$ ms per video) without complex modeling.

### v2

@@ Abstract
The rapid advancement of generative AI enables highly realistic synthetic videos, posing significant challenges for content authentication and raising urgent concerns about misuse. Existing detection methods often struggle with generalization and capturing subtle temporal inconsistencies. We propose ReStraV(Representation Straightening for Video), a novel approach to distinguish natural from AI-generated videos. Inspired by the “perceptual straightening” hypothesis [1, 2]—which suggests real-world video trajectories become more straight in neural representation domain—we analyze deviations from this expected geometric property. Using a pre-trained self-supervised vision transformer (DINOv2), we quantify the temporal curvature and stepwise distance in the model’s representation domain. We aggregate statistics of these measures for each video and train a classifier. Our analysis shows that AI-generated videos exhibit significantly different curvature and distance patterns compared to real videos. A lightweight classifier achieves state-of-the-art detection performance (e.g., 97.17% accuracy and 98.63% AUROC on the VidProM benchmark [3]), substantially outperforming existing image- and video-based methods. ReStraV is computationally efficient, offering a low-cost and effective detection solution. This work provides new insights into using neural representation geometry for AI-generated video detection.
Figure 1: The ReStraV method for AI-video detection. Inspired by “perceptual straightening,” our approach leverages the geometric insight that natural videos form “straighter” feature trajectories ($z_{i}$) than generated ones. The temporal curvature (Eq. 1) serves as the discriminative signal for detection.
@@ 1 Introduction
This work explores neural representational distance and curvature (formally defined in Eq.˜1) as discriminative signals for fake video detection. An overview of ReStraV is provided in Fig.˜1. According to the perceptual straightening hypothesis, natural inputs map to straight paths in neural representations while unnatural sequences form curved trajectories [1, 2]. This has been verified in neuroscience, psychophysics, on CNNs and LLMs [1, 23]. It is motivated by the idea that predictive coding might favor straight temporal trajectories in latent space because they are more predictable.
Taking inspiration, in this work, we hypothesize a distinction between natural and AI generated videos in artificial neural networks (ANNs). While ANNs may not perfectly replicate biological straightening [1, 2, 24], we expect their learned representations to show AI-generated videos as more curved in activation space than real videos. We surmise synthetic videos exhibit curvature patterns deviating from the lower curvature trajectories of real events, supported by differing ANN representational dynamics for natural versus artificial videos [24], as illustrated in Fig.˜2A and Fig.˜4.
To test this hypothesis, we use the DINOv2 ViT-S/14 pretrained visual encoder [25], chosen for its sensitivity to generative artifacts (Fig.˜2B). For each video, we extract frame-level complete set of patch embeddings and Classify token (CLS) from DINOv2’s final transformer block (block.11). From this trajectory, we quantify local curvature (angle between successive displacement vectors, measuring path bending) and stepwise distance (change magnitude between consecutive frame representations), as in Eq.˜1. We then derive descriptive statistics (mean, variance, min, max; examples in Fig.˜5) from these per-video time series of curvature and distance. These aggregated geometric features (Section˜5) are used by a lightweight classifier (Section˜6) to distinguish real from AI content.
ReStraV re-purposes DINOv2 as a “feature space” for temporal anomalies. DINOv2’s extensive training on natural data provides a latent space where real video trajectories should be characteristically smooth or “straight” (Fig.˜2A, Fig.˜4). Deviations, like increased jitter or erratic curvature often in AI videos, become discernible geometric signals of synthetic origin. Importantly, ReStraV is computationally efficient, processing videos in approximately 48 ms end-to-end (including DINOv2 forward pass). ReStraV is thus a low-cost alternative to resource-intensive methods (details in Section˜6). By exploiting ANN’s activation dynamics, ReStraV offers a simple, interpretable AI-video detection approach (experimental validation in Section˜7). Our contributions are as follows:
@@
Figure 2: (A) In pixel space (left), video trajectory metrics (curvature, distance; see Eq.˜1 for details) between natural vs. AI-generated videos show substantial overlap. In contrast, DINOv2 representations (right) straighten natural trajectories, clearly separating natural and AI-generated videos. (B) The mean curvature gap ($\Delta\theta$) between AI-generated and natural videos across various visual encoders. HVS-inspired models (red) exhibit negative deltas, straightening both natural and AI videos equally, while SSL models (green), particularly DINOv2, show the largest positive deltas.
@@ 2 Related work: detecting AI–generated videos
Detecting AI-generated video (see  Appendix˜A for an AI generative models overview) is becoming increasingly challenging. Many early detection efforts, including image-based detectors (CNNSpot [27], Fusing [28], Gram-Net [16], FreDect [29], GIA [30], LNP [31], DFD [32], UnivFD [33]), focused on spatial or frequency-domain artifacts within individual frames. However, their frame-centric nature limits their efficacy on videos, where temporal consistency is paramount.
Dedicated video detectors, such as adapted action recognition models (TSM [34], I3D [35], SlowFast [36]) and Transformer-based (X3D [37], MVIT-V2 [38], VideoSwin [39, 40], TPN [41], UniFormer-V2 [42], TimeSformer [43], DeMamba [44], aim to learn motion anomalies and temporal inconsistencies. While advancing temporal modeling, they require extensive training and may still struggle across rapidly evolving AI generative models. Those approaches may overlook a more fundamental signal: geometric distortions in the temporal trajectory of neural representations. We hypothesize that the geometric properties of these trajectories—reflecting the inherent smoothness and predictability of natural dynamics that generative models fail to replicate—offer a more robust signal for detection. Unlike related work in video quality assessment that also uses trajectories [45, 46], our focus is distinctly on detecting synthetic content, regardless of its perceptual quality.
@@ 3 Perceptual straightening definition
$\quad i=1,\dots,T-2$. The curvature $\theta_{i}$ (Equation 1), defined for each discrete step $i$ along the trajectory (ranging from 11 to T2T-2, where $T$ is the total number of sampled frames), is computed from the cosine similarity between successive displacement vectors $\Delta z_{i}$ and $\Delta z_{i+1}$. This geometric relationship is visualized in Figure 3. The figure depicts three consecutive frame embeddings ($z_{i},z_{i+1},z_{i+2}$) from the overall dashed trajectory. The orange vectors $\Delta z_{i}$ and $\Delta z_{i+1}$ represent the displacements between these embeddings, with their respective lengths being the stepwise distances $d_{i}$ and $d_{i+1}$. The green angle $\theta_{i}$ shows the turn at $z_{i+1}$. This angle, typically converted to degrees ($\theta_{i}^{\circ}=\theta_{i}\times\frac{180}{\pi}$), provides a measure of how sharply the representation trajectory bends at each step. These metrics, stepwise distance $d_{i}$ and curvature $\theta_{i}$, form the core of our geometric analysis.
@@ 4 Perceptual straightening of natural videos in DINOv2
The classic finding is that visual system model have lower curvature in the representation space compared to pixel space–this is called perceptual straightening [1, 2]. For our application, we want to compare the relative curvature of natural and AI generated videos in any representational space. Just for simplicity, we could call this real straightening as in real videos are less curved than AI videos. One might expect that a model with good perceptual straightening also has good real straightening.
To test this hypothesis, we analyze fourteen vision encoders across diverse families: Supervised CNNs (AlexNet [51], VGG-16 [52], ResNet-50 [53], and the texture-debiased SIN-ResNet-50 [54]); Self-Supervised (SimCLR-R50 [55], BYOL-R50 [56], CLIP [57], and DINOv2 [25]); Human Visual System-Inspired (a Gabor Filter Bank [58] and the LGN-V1 model [1]); Spatio-temporal (S3d [59], R(2+1)D [60], and MViT [61]); and an Adversarially Trained ResNet-50 [62]. Surprisingly, as shown in Fig.˜2B, we observe that the opposite seems to be the case: good perceptual straighteners actually make natural videos more curved than AI videos. HVS-inspired models achieve the strongest absolute straightening but do so indiscriminately for both real and AI videos, resulting in a negative curvature gap ($\Delta\theta<0$). In contrast, self-supervised models like DINOv2 reduce the curvature of natural videos, which align with their learned priors of real-world statistics, but do not regularize the trajectories of AI-generated videos, which violate these priors. This differential response creates a large, positive curvature gap ($\Delta\theta=45.46^{\circ}$ for DINOv2), which is the foundation of our method’s success. The negligible correlation between absolute straightening and detection capability ($\rho=-0.13,p=0.64$) confirms that detection performance hinges not on absolute straightening capability, but on differentially straightening natural versus synthetic videos. While artificial neural networks may not fully replicate the absolute perceptual straightening observed in the biological visual system [24], this relative effect is the key mechanism for our detection method.
This principle is clearly visualized in Fig.˜2A. Using matched pairs of real videos from Physics-IQ [26] and their AI-generated replicas, we see that geometric metrics overlap considerably in raw pixel space. However, in DINOv2’s representation space, the trajectories separate distinctly, offering a clean signal for detection. Further evidence is provided in Fig.˜4, which shows that this separation holds on a larger diverse dataset (VideoProM [3]). The t-SNE embeddings of curvature trajectories show natural videos (blue) forming a tight cluster, well-separated from the clusters of AI-generated videos (shades of red). This demonstrates that DINOv2’s features effectively surface temporal inconsistencies without any task-specific training. Refer to Section˜A.1 for trajectories samples.
@@
To implement this, we extract $T=24$ frames $\mathcal{Z}=(z_{1},\dots,z_{24})$ by sampling over a 22-second video duration. This 22-second window is suitable for the videos considered in Section˜7 ($\approx 2-5$s long with 123012-30 FPS), with an temporal based sample frame of $\Delta t=2\text{s}/(24-1)$. Nevertheless, we hypothesize that using longer videos could further enhance performance. Our choice of a $2s$ window with $24$ frames was found to provide an optimal trade-off between high detection accuracy and computational efficiency, as validated in our ablation studies (see B for details). Each $x_{i}$ is resized to $224\times 224$ pixels and normalized to $[0,1]$. These preprocessed frames are then encoded by the DINOv2 ViT-S/14 model [25]. The $384$ CLS (Classify) tokens and $196$ patch embedding (161616*16 patches of the 224224224*224 inputs) from its final transformer block (block.11). These token embeddings are then flattened and concatenated to form a single feature vector $z_{i}\in\mathbb{R}^{75648}$. The sequence of these vectors, $\mathcal{Z}$, forms the temporal trajectory in DINOv2’s representation space, from which temporal curvature and distance metrics are computed (Eq.˜1).
@@
Takeaway 1: By projecting videos into DINOv2’s representation space, geometric trajectory features (curvature & distance, Eq. 1) become indicators of synthetic origin, differentiating AI-generated videos from natural ones in a way that is not possible in raw pixel space.
@@ 5 Analyzing characteristics of perceptual trajectories
We select 50,000 AI-generated samples (10,000 each from Pika [47], VideoCraft2 [63], Text2Video-Zero [64], ModelScope [3], and Sora [65]) from VideoProM [3]. Concurrently, 50,000 natural videos are randomly chosen from DVSC2023 [48]. All videos are DINOv2 (ViT-S/14) encoded, and their aggregated statistical features are computed. Fig.˜5 illustrates these aggregated feature distributions. The top row shows distance features ($\mu_{d},\min_{d},\max_{d},\sigma_{d}^{2}$), characterizing inter-frame change magnitude and variability. The bottom row presents corresponding curvature features ($\mu_{\theta},\min_{\theta},\max_{\theta},\sigma_{\theta}^{2}$), reflecting angular changes between consecutive frame transitions.
@@ 6 Video classifier to detect AI-generated content
Given that DINOv2 representation distance $d$ and curvature $\theta$ differs significantly between natural and AI-generated videos, we evaluate if these features can be used in a lightweight, transparent, and easily replicated classifier without raw pixel processing or DINOv2 fine-tuning. We use the dataset from Section˜5 and apply a stratified 50/50 train/test split. Class priors are identical, and each subset is balanced among five AI models (Pika [47], VideoCraft2 [63], Text2Video-Zero [64], ModelScope [3] and Sora [65]). We sample frames and we obtain the signals of distance $\{d_{i}\}_{i=1}^{T-1}$ and curvature $\{\theta_{i}^{\circ}\}_{i=1}^{T-2}$ as detailed in Section 5. For classification, we construct a feature vector $y$ per video by combining direct signals and aggregated statistics from these trajectories. Specifically, $y$ concatenates seven distance values $d_{1},d_{2},\dots,d_{7}]$ and six curvature values: $\theta_{1}^{\circ},\theta_{2}^{\circ},\dots,\theta_{6}^{\circ}]$; and four statistical descriptors (mean, variance, minimum, maximum) for both $\{d_{i}\}$ and $\{\theta_{i}^{\circ}\}$. This results in a final feature vector $y\in\mathbb{R}^{21}$. To ensure our curvature-based features are detecting generative artifacts rather than hard scene cut frequency, we performed a robustness analysis detailed in Appendix˜C.
We consider only off-the-shelf models: logistic regression (LR), Gaussian Naive Bayes (GNB), random forest (RF; 400 trees, depth $\leq 6$), gradient boosting (GB; 200 rounds, learning rate 0.1), RBF-kernel SVM (calibrated by Platt scaling), and a two-layer MLP $(64\rightarrow 32)$. We perform no feature engineering or hyperparameter search beyond a 3-fold grid/random sweep. For each classifier, we optimize the decision threshold $\tau^{*}$ on the training set to maximize the F1-score. The chosen threshold $\tau^{*}$ was then applied unchanged to the test set. Inference cost is reported end-to-end ($\text{latency}=T_{\text{DINOv2}}+T_{\text{clf}}$), averaged over the test fold on a single NVIDIA RTX-2080 (see Appendix˜D). A DINOv2 forward pass (ViT-S/14, block 11, 8-frame batch) takes 43.6 ms. This constant is added to each classifier time ($T_{\text{clf}}$) in Fig.˜6.
Figure 6: Inference-time (msms) vs. accuracy (%) and AUROC (%) for ReStraV’s classifiers.
Table 1: Performance and inference time of ReStraV’s classifiers, balanced 50k/50k natural/AI-generated video test set from VideoProM [3], cf. Section˜5 for details. The best scores are bold, second best are underlined and the best method overall is highlighted in blue.
@@
Figure˜6 summarizes test performance. The MLP achieves the highest accuracy (94.17%), F1-score (96.48%), and AUROC (98.6%) (Fig.˜9A), followed by RF. Figure˜6 visualizes each classifier’s speed/accuracy. Models in the upper-left offer the best cost/benefit. The MLP (highlighted blue in Fig.˜6) achieves top AUROC and F1 while being within $\approx$2ms from GNB. The confusion matrices and ROC (Fig.˜9A and B in Section˜A.2) confirms low false positive/negative rates (cf. Section˜A.2 for decision boundaries visualization). Refer to Section˜A.3 for permutation feature importance analysis.
Takeaway 2: Simple, lightweight classifiers trained directly on geometric trajectory features achieve high classification accuracy ($\approx 94\%$) and AUROC ($\approx 99\%$), offering an efficient detection approach ($\approx 48$ ms per video) without complex modeling.